# Variational Low-Rank Adaptation for Personalized Impaired Speech Recognition

### v1

@@ Abstract
Speech impairments resulting from congenital disorders, such as cerebral palsy, down syndrome, or apert syndrome, as well as acquired brain injuries due to stroke, traumatic accidents, or tumors, present major challenges to automatic speech recognition (ASR) systems. Despite recent advancements, state-of-the-art ASR models like Whisper still struggle with non-normative speech due to limited training data availability and high acoustic variability. Moreover, collecting and annotating non-normative speech is burdensome: speaking is effortful for many affected individuals, while laborious annotation often requires caregivers familiar with the speaker. This work introduces a novel ASR personalization method based on Bayesian Low-rank Adaptation for data-efficient fine-tuning. We validate our method on the English UA-Speech dataset and a newly collected German speech dataset, BF-Sprache, from a child with structural speech impairment. The dataset and approach are designed to reflect the challenges of low-resource settings that include individuals with speech impairments. Our method significantly improves ASR accuracy for impaired speech while maintaining data and annotation efficiency, offering a practical path toward inclusive ASR.
@@ 1 Introduction
These challenges are further aggravated for non-English languages due to the lack of representative data, limited linguistic tools, and the need for language-specific adaptation strategies [5, 7]. German remains under-resourced for non-normative speech, particularly for childrenn, with limited publicly available datasets [8, 9, 10].
@@
Bayesian Neural Network (BNN)-based approaches offer promising alternatives for improving ASR robustness in low-data, high-variability settings [16], and have been used to maintain robustness across continuous adaptation [17]. BNN approaches may provide key benefits in bridging parameter efficient fine-tuning methods to the high variability data of impaired speech. To this end, our work proposes:
@@
    Variational Low-rank Adaptation framework. We introduce a principled Bayesian low-rank adaptation method called VI LoRA that captures uncertainty during fine-tuning. This enables robust personalization with significantly less data, while maintaining parameter efficiency, crucial for modeling speech with high acoustic variability.
@@
    Data-driven prior estimation. We develop a prior estimation approach that better captures the multi-modal distribution of layer-wise weight variations in state-of-the-art ASR architectures. We show that this approach improves robustness and stability during training.
@@ 2.2 Bayesian Low-rank Adaptation Framework
To address the challenges of low data availability, overfitting, and overparameterization encountered when fine-tuning large models [12], we propose a Bayesian low-rank adaptation framework. This approach enhances the standard low-rank adaptation (LoRA) [13], a prominent parameter-efficient fine-tuning (PEFT) technique [19]. LoRA adapts a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ by freezing $W_{0}$ and introducing a trainable low-rank update $\Delta W_{0}$. This update is parameterized as the product of two smaller matrices, $B\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $A\in\mathbb{R}^{r\times d_{\mathrm{in}}}$, where the rank $r\ll\min(d_{\mathrm{in}},d_{\mathrm{out}})$. The adapted weight matrix then becomes $W_{0}+\frac{\alpha}{r}BA$, with $\alpha$ being a scaling factor.
@@
here, $p(A,B)$ is the prior distribution over the LoRA matrices. The term $\mathbb{E}_{q_{\phi}(A,B)}[\log p(\mathcal{D}|A,B)]$ is the expected log-likelihood of the training data, corresponding to the task-specific loss (e.g., cross-entropy for ASR) computed with samples from $q_{\phi}(A,B)$, typically estimated using Monte Carlo sampling. Let $N$ be the total number of LoRA layers where VI is applied. The KL divergence term in the ELBO, $\mathrm{KL}[q_{\phi}(A,B)||p(A,B)]$, is theoretically a sum over all $N$ layers. However, numerical instabilities, especially in early training stages before posteriors stabilize, can lead to non-finite KL values for some layers. To ensure robust optimization, our final loss function $\mathcal{L}_{\mathrm{VI}}$ employs an average over layers yielding finite (non-NaN, non-Inf) KL divergence terms in the current optimization step:
@@
where $\mathcal{W}^{(l)}$ represents the set of all weights in the original matrix $W_{0}^{(l)}$, and $\overline{w}^{(l)}$ is their mean. Our empirical analysis of these layer-wise standard deviations, $\{\hat{\sigma}_{p}^{(l)}\}_{l=1}^{N}$ (across $N=288$ target layers in our experiments), reveals a distinct bimodal distribution, as illustrated in Figure 1. A simple Gaussian Mixture Model was used to find the optimal $\mu$ for the layer specific $\sigma$ prior.
@@ 3.1 Datasets
The English UA-Speech dataset [24] is widely used benchmarks in dysarthric speech recognition research and has laid the foundation for numerous deep learning-based works [26, 27]. The dataset consists of recordings from 19 speakers with varying degrees of dysarthria, along with recordings from 10 control speakers. The dataset emphasizes isolated word-level utterances, comprising 455 distinct words including digits, letters, and phonetically rich uncommon words.
@@
To measure the forgetting of the already learned distribution of normal speech we used a split in the respective language (German for BF-Sprache and English for UA-Speech) of the widely known Mozilla Common Voice Dataset.
@@ 3.2 Model Evaluation
Table 1: Results of different models on normative and non-normative speech. WD refers to Weight Decay, DP to Dual Prior, SP to Single Prior, KL to 10% KL[q||p][q||p].
@@ 4 Results and Analysis
Table 2: Results on the UA-Speech dataset for different adaptation methods, including a 0-shot baseline. All methods are evaluated on both normative and non-normative speech.
Setup Speech Type CER WER
@@
Table 3: CER and WER on non-normative speech for different training dataset sizes and adaptation methods.
@@
Table 1 presents a comparative analysis of different adaptation strategies, including standard LoRA, full parameter fine-tuning, and VI LoRA with and without a KL regularization term. All models were fine-tuned on BF-Sprache. The results indicate that VI LoRA, when regularized with a 10% KL[q||p][q||p] term, demonstrates a compelling trade-off. This configuration achieves the lowest CER and WER on the target non-normative speech. Concurrently, it exhibits the least forgetting of normative speech, as evidenced by its leading CER and WER scores, surpassing both standard LoRA and full parameter fine-tuning which show higher error rates on normative data. Those findings translate well onto multiple english speaker of diffferent intelligibility levels from the UA-Speech dataset as shown in Table 2.
We speculate that the performance of VI LoRA with KL divergence regularization stems from its ability to effectively adapt to non-normative speech characteristics while mitigating catastrophic forgetting. The KL divergence term acts as a regularizer, penalizing significant deviations of the adapted LoRA weights ($q$) from the original pre-trained weight distribution ($p$). This constraint likely encourages the model to learn the specific variations of non-normative speech more parsimoniously, preventing overly aggressive updates that could shift the weights too far from their initial state beneficial for normative speech. This allows for controllable adaptation without sacrificing much of the model’s generalization ability to the pre-trained normative speech patterns.
Table 3 reports CER/WER on non-normative speech across varying training set sizes. VI LoRA consistently outperforms all baselines, with its advantage being most pronounced when less data is available. While full fine-tuning underperforms even standard LoRA in low-data settings, it surpasses LoRA once the full training set is used. This trend is particularly evident in WER.
As illustrated in Table 4, a qualitative analysis of the transcription outputs reveals a critical distinction between the error patterns of full fine-tuning and our proposed VI LoRA model on out-of-distribution (OOD) phrases. The fully fine-tuned model consistently exhibits a form of structured hallucination. For instance, it transcribes the Japanese place name "Higashirinkan" as the grammatically plausible but semantically unrelated German sentence "Ein Gassi rennt da." This suggests the model defaults to familiar linguistic patterns when faced with novel acoustic signals, effectively pattern-matching to the closest structure in its learned distribution.
@@
Table 4: Comparison of transcription examples for out-of-distribution phrases. Phonetic transcriptions (IPA) are provided in italics to aid interpretation for non-German speakers.
@@ 5 Discussion and Conclusions
In this work, we propose a novel personalization framework using Bayesian LoRA. Our experimental results demonstrate substantial improvements in recognition accuracy for speech-impaired individuals across different intelligibility levels and languages, while maintaining efficient parameter usage.
Our experiments revealed several key insights. Our dual prior approach to VI LoRA (Table 1) showed particularly strong performance, reducing non-normative speech CER to 20.09% (compared to 21.33% for single prior) while maintaining reasonable normative speech performance. This suggests that modeling the bimodal distribution of pre-trained weights significantly improves adaptation capability.
The cross-lingual results are especially encouraging, with our framework generalizing effectively from English to German despite the latter’s phonological complexity. Furthermore, the observed trade-off between performance on non-normative and normative speech highlights a key design consideration: systems optimized solely for impaired speech may sacrifice generalization, suggesting that multi-objective training strategies could further improve generalizability.
While promising, our work has some limitations. Our pipeline assumes independent factorization of $q_{\phi}(A,B)$ by disentangling it as $q_{\phi_{A}}(A)q_{\phi_{B}}(B)$. While computationally efficient, it may not best capture the interactions between LoRA adapter matrices [28, 29]. Also, educational considerations must be noted, as systems that rely heavily on contextual pattern matching like Whisper may discourage children from learning correct grammar, emphasizing the importance of accurate word-by-word transcription in educational contexts. The main limitation of this study however, remains the small speaker pool in the BF-Sprache dataset, due to previous resource and ethical approval constraints. With the ethical approval now secured, our immediate future work will focus on expanding this dataset by recruiting a larger, more diverse group of speakers across various conditions and intelligibility levels.
Our further work will focus on expanding the speaker-base in BF-Sprache over different conditions and intelligibility-levels as well as incorporating VI LoRA in a active learning setting for continuous speaker specific adaption, leveraging its robustness against forgetting of normative speech.

### v2

@@ Abstract
Speech impairments resulting from congenital disorders, such as cerebral palsy, Down syndrome, or Apert syndrome, as well as acquired brain injuries due to stroke, traumatic accidents, or tumors, present major challenges to automatic speech recognition (ASR) systems. Despite recent advancements, state-of-the-art ASR models like Whisper still struggle with non-normative speech due to limited training data availability and high acoustic variability. Moreover, collecting and annotating non-normative speech is burdensome: speaking is effortful for many affected individuals, while laborious annotation often requires caregivers familiar with the speaker. This work introduces a novel ASR personalization method based on Bayesian Low-rank Adaptation for data-efficient fine-tuning. We validate our method on the English UA-Speech dataset and a newly collected German speech dataset, BF-Sprache, from an individual with structural speech impairment. Both the dataset and the approach are designed to reflect the challenges of low-resource settings that include individuals with speech impairments. Our method significantly improves ASR accuracy for impaired speech while maintaining data and annotation efficiency, offering a practical path toward inclusive ASR.
@@ 1 Introduction
These challenges are further aggravated for non-English languages due to the lack of representative data, limited linguistic tools, and the need for language-specific adaptation strategies [2, 30]. German remains under-resourced for non-normative speech, particularly for children, with limited publicly available datasets [25, 8, 7].
@@
Bayesian Neural Network (BNN)-based approaches offer promising alternatives for improving ASR robustness in low-data, high-variability settings [18], and have been used to maintain robustness across continuous adaptation [6]. BNN approaches may provide key benefits in bridging parameter-efficient fine-tuning methods to the high variability data of impaired speech. Prior works primarily leverage Bayesian low-rank adaptation (LoRA) for efficiency, either via capacity-reducing pruning and quantization [5, 17] or static post-hoc analysis [32, 28], which risks underfitting the complex acoustic features of disordered speech. Conversely, our method employs variational inference (VI) as a dynamic training regularizer that retains full capacity while elastically constraining adaptation to the pre-trained weight structure. To this end, our work proposes:
@@
    Variational Low-rank Adaptation framework. We introduce a Bayesian LoRA method called VI LoRA that captures uncertainty during fine-tuning. This enables robust personalization with significantly less data, while maintaining parameter efficiency, crucial for modeling speech with high acoustic variability.
@@
    Data-driven prior estimation. We develop a prior estimation approach that better captures the multi-modal distribution of layer-wise weight variations in state-of-the-art ASR architectures.
@@ 2.2 Bayesian Low-rank Adaptation Framework
Our approach addresses the challenges of low data availability, overfitting, and over-parameterization when fine-tuning large models [22] and enhances the standard low-rank adaptation (LoRA) [10], a prominent parameter-efficient fine-tuning (PEFT) technique [14]. LoRA adapts a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ by freezing $W_{0}$ and introducing a trainable low-rank update $\Delta W_{0}$. This update is parameterized as the product of two smaller matrices, $B\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $A\in\mathbb{R}^{r\times d_{\mathrm{in}}}$, where the rank $r\ll\min(d_{\mathrm{in}},d_{\mathrm{out}})$. The adapted weight matrix then becomes $W_{0}+\frac{\alpha}{r}BA$, with $\alpha$ being a scaling factor.
@@
here, $p(A,B)$ is the prior distribution over the LoRA matrices. The term $\mathbb{E}_{q_{\phi}(A,B)}[\log p(\mathcal{D}|A,B)]$ is the expected log-likelihood of the training data, corresponding to the task-specific loss (e.g., cross-entropy for ASR) computed with samples from $q_{\phi}(A,B)$, typically estimated using Monte Carlo sampling. Let $N$ be the total number of LoRA layers where VI is applied. The KL divergence term in the ELBO, $\mathrm{KL}[q_{\phi}(A,B)||p(A,B)]$, is theoretically a sum over all $N$ layers. However, numerical instabilities, especially in early training stages before posteriors stabilize, can lead to non-finite KL values for some layers. To ensure robust optimization, our final loss function $\mathcal{L}_{\mathrm{VI}}$ employs an average over layers yielding a finite (non-NaN, non-Inf) KL divergence terms in the current optimization step:
@@
where $\mathcal{W}^{(l)}$ represents the set of all weights in the original matrix $W_{0}^{(l)}$, and $\overline{w}^{(l)}$ is their mean. Our empirical analysis of these layer-wise standard deviations, $\{\hat{\sigma}_{p}^{(l)}\}_{l=1}^{N}$ (across $N=288$ target layers in our experiments), reveals a distinct bimodal distribution, as illustrated in Figure 1. A simple Gaussian Mixture Model was used to find the optimal $\mu$ for the layer-specific $\sigma$ prior.
@@ 3.1 Datasets
The English UA-Speech dataset [13] is widely used in dysarthric speech recognition research and has laid the foundation for numerous deep learning-based works [27, 4]. It consists of recordings from 19 speakers with varying degrees of dysarthria, along with recordings from 13 control speakers. We exclusively used the $\approx 66h$ of dysarthic speech. The dataset emphasizes isolated word-level utterances, comprising 455 unique words including digits, letters, and phonetically rich uncommon words.
@@
To measure the forgetting of the already learned distribution of normal speech we used a split in the respective language (German or English) of the widely known Mozilla Common Voice Dataset [1] for validation and testing.
@@ 3.2 Model Evaluation
Table 1: Results of different models on non-normative (BF-Sprache) and normative (CommonVoice) speech. WD refers to Weight Decay, DP to Dual Prior, SP to Single Prior, KL to 10% KL[q||p][q||p].
@@ 4 Results and Analysis
Table 2: Results on the UA-Speech (non-normative) and CommonVoice (normative) dataset for different adaptation methods, including a 0-shot baseline, relative to full fine-tuning (= 100%).
Setup Speech Type rel. CER rel. WER
@@
Table 3: CER and WER on BF-Sprache (100% $\approx$ 2h) for different sizes of the training set and adaptation methods.
@@
Table 1 presents a comparative analysis of different adaptation strategies, including standard LoRA, full parameter fine-tuning, and VI LoRA with and without a KL regularization term. All models were fine-tuned on BF-Sprache. The results indicate that VI LoRA, when regularized with a 10% KL[q||p][q||p] term, demonstrates a compelling trade-off. This configuration achieves the lowest CER and WER on the target non-normative speech. Concurrently, it exhibits the least forgetting of normative speech, as evidenced by its leading CER and WER scores, surpassing both standard LoRA and full parameter fine-tuning, which show higher error rates on normative data. These findings generalize across English speakers with varying intelligibility levels in the UA-Speech dataset (Table 2). The table reports relative performance differences from full fine-tuning, aggregated over all dysarthric speakers.
We speculate that the performance of VI LoRA with KL divergence regularization stems from its ability to effectively adapt to non-normative speech characteristics while mitigating catastrophic forgetting. The KL divergence term acts as a regularizer, penalizing significant deviations of the adapted LoRA weights ($q$) from the original pre-trained weight distribution ($p$). This constraint likely encourages the model to learn the specific variations of non-normative speech more parsimoniously, preventing overly aggressive updates that could shift the weights too far from their initial state, which is beneficial for normative speech. This allows for controllable adaptation without sacrificing much of the model’s generalization ability to the pre-trained normative speech patterns.
Table 3 reports CER/WER on non-normative speech in the BF-Sprache dataset across varying training set sizes. The test set is fixed, while training sets are constructed from randomly selected instances of the remaining data. VI LoRA consistently outperforms all baselines, with its advantage being most pronounced when less data is available. While full fine-tuning underperforms even standard LoRA in low-data settings, it surpasses LoRA once the full training set is used. This trend is particularly evident in WER.
As illustrated in Table 4, a qualitative analysis of the transcription outputs reveals a critical distinction between the error patterns of full fine-tuning and our proposed VI LoRA model on out-of-distribution (OOD) phrases, specifically targeting semantically rare or hyper-locally relevant terms (e.g., “Wiedikon”) that lack strong language model priors. The fully fine-tuned model consistently exhibits a form of structured hallucination. For instance, it transcribes the Japanese place name "Higashirinkan" as the grammatically plausible but semantically unrelated German sentence "Ein Gassi rennt da." This suggests the model defaults to familiar linguistic patterns when faced with novel acoustic signals, effectively pattern-matching to the closest structure in its learned distribution.
@@
Table 4: Qualitative comparison of transcription examples for out-of-distribution phrases. Phonetic transcriptions (IPA) are provided in italics to aid interpretation for non-German speakers.
@@ 5 Discussion and Conclusions
In this work, we propose a novel personalization framework using Bayesian LoRA. Our experimental results demonstrate substantial improvements in recognition accuracy for speech-impaired individuals across different intelligibility levels and languages, while maintaining efficient parameter usage. Our experiments revealed several key insights. Our dual prior approach to VI LoRA (Table 1 for BF-Sprache and Table 2 for UA-Speech) showed particularly strong performance, reducing non-normative speech CER to 20.09% (compared to 21.33% for single prior) while maintaining reasonable normative speech performance. This suggests that modeling the bimodal distribution of pre-trained weights significantly improves adaptation capability. The cross-lingual results are especially encouraging, with our framework generalizing effectively from English to German despite the latter’s phonological complexity. Furthermore, the observed trade-off between performance on non-normative and normative speech highlights a key design consideration: systems optimized solely for impaired speech may sacrifice generalization, suggesting that multi-objective training strategies could further improve generalizability. While promising, our work has some limitations. Our pipeline assumes independent factorization of $q_{\phi}(A,B)$ by disentangling it as $q_{\phi_{A}}(A)q_{\phi_{B}}(B)$. While computationally efficient, it may not best capture the interactions between LoRA adapter matrices [33, 16]. The main limitation remains the small speaker pool in the BF-Sprache dataset, due to previous resource and ethical approval constraints. With the ethical approval now secured, our immediate future work will focus on expanding this dataset by recruiting a larger, more diverse group of speakers across various conditions and intelligibility levels. Our further work will focus on expanding the speaker base in BF-Sprache over different conditions and intelligibility levels as well as incorporating VI LoRA in an active learning setting for continuous speaker-specific adaptation.